OpenStudy (jozelynw):

2. Then, explain to Iris how the key features of this local snail population compares to the key features of the invasive population. Where it says then, that is the other part of the question. I need help on that part.

OpenStudy (anonymous):

hmm...

OpenStudy (jozelynw):

what do you think @MAli13chineta

OpenStudy (anonymous):

i dont really understand your question is there also a picture?

OpenStudy (jozelynw):

no it's a project in math with four questions. I don't understand the second part in the question, i already did the exponent part of the question. this is the whole question 2. A local snail population grows according to the function g(x) = 200(1.03)2x. 2.)Demonstrate the steps to convert g(x) into an equivalent function with only x as the exponent. Then, explain to Iris how the key features of this local snail population compares to the key features of the invasive population.

OpenStudy (jozelynw):

@MAli13chineta @mathstudent55 @mathmate

OpenStudy (jozelynw):

@mathmate what do you think?

OpenStudy (mathmate):

Do you have a problem with the conversion part? Use the identity in the law of exponents: \(\large a^{bx}=a^b\times a^x\)

OpenStudy (jozelynw):

No the answer is g(x)=200(1.0609)^x @mathmate . I'm talking about the second question the question asks